OpenStudy (anonymous):

log4(5x + 1)=2 solve

OpenStudy (anonymous):

same as previous one \[\log_{a}b=c\] then \[b=a^c\] substitute values of a,b,c from the given expression

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

4^2 = 5x + 1 16 = 5x + 1 15 = 5x 3 = x